OpenStudy (anonymous):

In a 45°-45°- 90° triangle, if the legs each have a length of 4, then what is the length of the hypotenuse?

OpenStudy (mertsj):

You could use the Pythagorean Theorem

OpenStudy (anonymous):

Pythag?

OpenStudy (mertsj):

Or if you know your special right triangles, you could use that.

OpenStudy (anonymous):

4^2+4^2=C^2

OpenStudy (anonymous):

im not really sure what im suppose to do.

OpenStudy (mertsj):

Do you know the pythagorean theorem?

OpenStudy (anonymous):

not really explain please? im having trouble with this

OpenStudy (mertsj):



OpenStudy (mertsj):

The Pythagorean Theorem is for right triangles. It says to square each leg, and those answers together and that is the square of the hypotenuse.

OpenStudy (anonymous):

A^2+b^2=C^2

OpenStudy (mertsj):